OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax and x represents the original price of the item: C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale? Remember, do not include the $ symbol.

Vocaloid (vocaloid):

all you need to do is let x = 100 and calculate 0.765x + 0.06(0.765x)

OpenStudy (anonymous):

wait what i dont get that

Vocaloid (vocaloid):

x = 100 0.765*x + 0.06*(0.765)*x = ___ fill in the blank

OpenStudy (anonymous):

wait whats with the parenthasis

Vocaloid (vocaloid):

it's part of the original equation... please read the problem one more time

OpenStudy (anonymous):

do i multiply the 0.765 to the 100

Vocaloid (vocaloid):

yes

OpenStudy (anonymous):

i got 76.5

Vocaloid (vocaloid):

right, keep going with the rest of the equation

OpenStudy (anonymous):

i add 76.5 to 0.06

Vocaloid (vocaloid):

nope, remember order of operations. multiply first, then add

Vocaloid (vocaloid):

(0.765*x) + (0.06*(0.765)*x) = 76.5 + 0.06*(0.765)*x = ?

OpenStudy (anonymous):

76.56

Vocaloid (vocaloid):

that isn't the correct answer, please try one more time we can split the problem into two parts: 1. 0.765*x 2. 0.06*(0.765)*x you found the first part, which is 76.5 now find the second part: 0.06*(0.765)*x = ?

OpenStudy (anonymous):

0.0459

Vocaloid (vocaloid):

not quite, try one more time... remember that x = 100 0.06*(0.765)*x = ?

OpenStudy (anonymous):

4.59

Vocaloid (vocaloid):

right, now we add parts 1 and 2 together 76.5 + 4.59 = ?

OpenStudy (anonymous):

81.09

Vocaloid (vocaloid):

yup, that's our final answer